The moment of inertia of slim rod about a transverse axis through one
end is M(L^2)/3, where M is the mass of the rod and L its length, and L^2 is L
squared. What is the moment of inertia of the same rod if we choose a new
parallel axis at a distance of L/3 from the original axis?
1
Expert's answer
2020-12-04T09:34:52-0500

From the Huygens–Steiner theorem:

"J=J_C+Md^2,"

"J_C=\frac{ML^2}{12},"

"d=\frac L2 -\frac L3=\frac L6,"

so "J=\frac{ML^2}{12}+\frac{ML^2}{36}=\frac{ML^2}{9}."